OpenStudy (anonymous):

If f(X) = e^x, which of the following lines is an asymptote to the graph of f? a) Y = 0 b) x=0 c) y=x d) y=-x e) y = 1

OpenStudy (anonymous):

y = 0

OpenStudy (anonymous):

How did you get to that?? Is it because it runs along the zero, why wasnt it E?

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (anonymous):

You would check the end behavior of the graph by evaluating the limits at infinity: \[\lim_{x\to\infty}e^x=\cdots~~~~~~~~~~\text{and}~~~~~~~~~~~\lim_{x\to-\infty}e^x=\cdots\]